OpenStudy (anonymous):

There is a parabola A: y= x^2 - 4x - 5 The equation of parabola which moved the parabola A symmetrically with the origin is y= ?

OpenStudy (anonymous):

y=x^2 - 4x - 5 y=x^2 - 4x +4 -9 y=(x-2)^2 -9 y=9=(x-2)^2

OpenStudy (anonymous):

y+9=(x-2)^2

OpenStudy (anonymous):

Is that the answer? Is the solution for this problem is to find something like the invers of y?

OpenStudy (anonymous):

no this i sthe original parabola..

OpenStudy (anonymous):

actually i am not clear what the question mean..

OpenStudy (anonymous):

can you post complete and the original problem...

OpenStudy (anonymous):

It's the question number 2. Also if it's okay, can you tell me how to answer number 3 part 2 (the area of portion)?

OpenStudy (anonymous):

for y+9=(x-2)^2 the vertex is given by (2,-9)

OpenStudy (anonymous):

Ah, yes, I forgot to put the -9 there.

OpenStudy (anonymous):

you have written it may be they mean to say y = x^2 (i am not sure abt this)

OpenStudy (anonymous):

For Area you need to know integration..

OpenStudy (anonymous):

Ooh, thanks. Finally got the correct answer for number 3.